Answer the two questions and show complete solution.

2. Find two unit vectors that are parallel to the
1
Expert's answer
2020-04-20T11:47:43-0400

Let the line be y=3x+4

Let's find two unit vectors that are parallel to that line.

The slope the given line is 3.

Suppose that, this line makes an angle of "\theta" with the +ve direction of the X−Axis.

Then, the unit vector "\overline{u}=(\cos\theta,\sin\theta)" is parallel to the line.

By the definition of slope, we have

"\tan\theta=3"

therefore

"\sec^2\theta=1+\tan^2\theta=10"

"cos\theta=\frac{1}{\sec\theta}=\pm\frac{1}{\sqrt{10}}"

"\sin\theta=\sqrt{1-cos^2\theta}=\sqrt{1-\frac{1}{10}}=\pm\frac{3}{\sqrt{10}}"

Hence "\overline{u}(\frac{1}{\sqrt{10}},\frac{3}{\sqrt{10}})" and "\overline{v}(-\frac{1}{\sqrt{10}},-\frac{3}{\sqrt{10}})" are two vectors, parallel to the line y=3x+4